OpenStudy (anonymous):

Train A & B are traveling in the same direction on parallel tracks. Train A is traveling at 40 mph & Train B at 48 mph. Train A passes a station at 2:20 am. If Train B passes the same station at 2:35 am, what time will Train A catch up to Train A?

OpenStudy (anonymous):

OOps, I meant train B catch up to train A

OpenStudy (watchmath):

Train A leaves 15 minutes = 1/4 hour earlier than train B (from the station) After x hours the distance traveled by both train are Distance train A= (1/4)(40)+40x Distance train B= 48x Train B catch up train A if the two distance are equal \(10+40x=48x\) \(8x=10\) \(x=5/4\) hour = 1 hour 15 minutes So Train B catcch up Train A at 3:50 am

OpenStudy (anonymous):

Thank you!

OpenStudy (watchmath):

If my answer help you please press the good answer button :D :D